# Information About Other Agents in Mechanism Design

### v3

@@ 3 Partial implementation and the direct mechanism
In this section we show that providing additional information to agents about the types of other agents is not helpful for the planner interested only in the partial implementation of an SCF. In our framework, this means that if $F$ is not implementable with respect to an uninformative signal structure, then it is not implementable with respect to any signal structure $\{\sigma_{i}\}_{i\in I}$. For the remainder of this paper, write $\phi_{i}$ to mean an uninformative signal about $\Theta_{-i}$ for each $i.$
@@
The law of total expectation implies that  (3) is just a weighted average of the equations  (3) for $1\leq\ell\leq k.$ This implies the following proposition.
@@
It is well-known that the direct mechanism generally has undesired equilibria. The mechanism exhibited by Jackson (1991) to fully implement an SCF $F$ allows each agent to report, among other information, an alternate SCF that can be used to signal a deception to the planner, provided $F$ is Bayesian monotonic. In this paper, it will be useful to keep track of which agents have an incentive to alert the planner of a given deception, so we introduce the following notion.
@@
The inequalities are exactly the ones from the definition of Bayesian monotonicity. Since a deception is equivalent to a strategy profile in the direct mechanism, the only deceptions $\alpha$ such that $F\circ\alpha\neq F$ that need to be undermined in order to guarantee the full implementation of $F$ are (untruthful) equilibria of the direct mechanism. For any incentive-compatible SCF $F$, write $\mathcal{A}_{F}$ to be the set of deceptions $\alpha$ such that $F\circ\alpha\neq F$ and constitute an equilibrium of the direct mechanism when agents do not have additional information about other agents. Let $\mathcal{A}_{F}^{u}\subset\mathcal{A}_{F}$ be the set of deceptions that can be undermined by some agent $i,$ again when agents have no additional information about other agents. Then we have the following.
@@
Let $F$ be an incentive-compatible SCF. In economic environments, $F$ is fully implementable if and only if $\mathcal{A}_{F}^{u}=\mathcal{A}_{F}$.
The lemma is a direct implication of the full Bayesian implementation result of Jackson (1991) but will prove useful in our general results.
4 Main example
In this section we exhibit an SCF that is not fully implementable with respect to $\{\phi_{i}\}_{i\in I}$ but is fully implementable with respect to some informative signal structure $\{\sigma_{i}\}_{i\in I}.$ The key constraint considered in this section is Bayesian monotonicity. In particular, we show how a signal about the types of other agents can incentivize an agent to alert the planner of a deception, which the agent would not have an incentive to do without this information.
The environment has 3 agents, who each have two types, drawn non-independently from a full-support common prior over the eight possible agent type profiles. Under the planner’s desired SCF, which we will first show is incentive-compatible, there are two agents who each have an incentive to misreport her type given that the other is as well. Therefore, there is an equilibrium of the direct mechanism in which these two agents are simultaneously deceiving. Next we show that this equilibrium arises from a deception that violates Bayesian monotonicity, implying that any mechanism that partially implements the SCF will have this undesirable equilibrium. Thus the SCF is not fully implementable when agents have no additional information about other agents’ types. However, there is a third agent for which truth-telling is a dominant strategy. We show that increasing the information she has about the types of the other two agents incentivizes her to alert the planner about this deception, implying that the SCF can satisfy Bayesian monotonicity under a more informative information structure. Finally, we show that this modification to the third agent’s information allows the SCF to become fully implementable.
We now present the example formally, which will illustrate the main points of the general result. Suppose the environment consists of three agents $i=1,2,3$, where each agent $i$ has types $\theta_{i},\theta_{i}^{\prime}.$ There are 6 possible outcomes $A=\{a,b,c,d,e,f\}.$ The SCF $F$ is as follows.
$F\mid\theta_{3}$ $\theta_{2}$ $\theta_{2}^{\prime}$
$\theta_{1}$ $a$ $b$
$\theta_{1}^{\prime}$ $e$ $d$
$F\mid\theta_{3}^{\prime}$ $\theta_{2}$ $\theta_{2}^{\prime}$
$\theta_{1}$ $a$ $c$
$\theta_{1}^{\prime}$ $e$ $d$
The agents have the following preferences over $A$.
Type Preferences
$\theta_{1}$ $e\,(5.5)\succ a\,(4)\succ b,c\,(3)\succ d\,(2)\succ f\,(1)$
$\theta_{1}^{\prime}$ $b,c\,(5.5)\succ d\,(4)\succ e\,(3)\succ a\,(2)\succ f\,(1)$
$\theta_{2}$ $f\,(15)\succ b,c\,(4.5)\succ a\,(3)\succ e\,(2)\succ d\,(1)$
$\theta_{2}^{\prime}$ $f\,(15)\succ e\,(4.5)\succ d\,(3)\succ b,c\,(2)\succ a\,(1)$
$\theta_{3}$ $a\,(5)\succ b\,(4)\succ e\,(3)\succ d\,(2)\succ c,f\,(1)$
$\theta_{3}^{\prime}$ $a\,(5)\succ c\,(4)\succ e\,(3)\succ d\,(2)\succ b,f\,(1)$
Each agent is indifferent between two outcomes and has strict preferences over the others. The numbers in parentheses represent the utilities of each type of each agent from the outcomes.
Let $\mu$ be the common prior on $\Theta$ that places $1/6$ probability on each of the type profiles $(\theta_{1},\theta_{2},\cdot)$, $(\theta_{1}^{\prime},\theta_{2}^{\prime},\cdot),$ and places $1/3$ probability on each of the type profiles $(\theta_{1}^{\prime},\theta_{2},\cdot)$ and $(\theta_{1},\theta_{2}^{\prime},\cdot).$ Independently, each of the type profiles $(\cdot,\cdot,\theta_{3}),(\cdot,\cdot,\theta_{3}^{\prime})$ has a $1/2$ probability. In other words, the type of agent 3 is independent of the types of agents 1 and 2, but type $\theta_{1}$ of agent 1 infers $1:2$ odds of agent 2 being type $\theta_{2}$ vs. type $\theta_{2}^{\prime}$, while type $\theta_{1}^{\prime}$ of agent 1 infers $2:1$ odds of agent 2 being type $\theta_{2}$ vs. type $\theta_{2}^{\prime}$.
The outcome $f$ exists to make this environment economic, so incentive compatibility and Bayesian monotonicity are necessary and sufficient for full implementation.
The claim is that $F$ is partially but not fully implementable with respect to $\{\tau_{i}\}$ if $\tau_{i}$ is the uninformative signal for all $i.$ In particular $F$ is incentive-compatible, but there is a deception that creates a violation of Bayesian monotonicity, meaning that it cannot be ruled out by any mechanism.
4.1 Incentive compatibility
To see that $F$ is incentive-compatible, truth-telling must be an equilibrium of the direct mechanism. For agent 1, observe
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1})\mid\theta_{1})=\frac{1}{3}u_{1}(a,\theta_{1})+\frac{1}{3}u_{1}(b,\theta_{1})+\frac{1}{3}u_{1}(c,\theta_{1})=10/3$
$\geq\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}^{\prime}),\theta_{1})\mid\theta_{1})=\frac{1}{3}u_{1}(e,\theta_{1})+\frac{2}{3}u_{1}(d,\theta_{1})=19/6.$
Similarly,
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}^{\prime}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=10/3\geq\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=19/6.$
Therefore, the incentive compatibility constraint holds for agent 1. A similar argument shows the incentive compatibility constraint holds for agent 2. For agent 3, truth-telling is a dominant strategy, as type $\theta_{3}$ strictly prefers outcome $b$ to $c$ while type $\theta_{3}^{\prime}$ strictly prefers outcome $c$ to $b.$ Thus, $F$ is incentive-compatible.
4.2 Bayesian monotonicity
To show that $F$ is partially but not fully implementable, we need to exhibit a deception $\alpha$ such that $F\circ\alpha\neq F$ that violates Bayesian monotonicty. We claim that $\alpha=((\theta_{1}^{\prime},\theta_{1}),(\theta_{2}^{\prime},\theta_{2}),(\theta_{3},\theta_{3}^{\prime}))$ creates a violation of Bayesian monotonicity. In order for this to be true, it must be an equilibrium of the direct mechanism but fail the Bayesian monotonicity constraint described above. To see that $\alpha$ is an undesired equilibrium of the direct mechanism, suppose agent 1 deviates to $(\theta_{1}^{\prime},\theta_{1}).$ Then type $\theta_{2}$ of agent 2 gets an expected payoff of
$\frac{2}{3}u_{2}(a,\theta_{2})+\frac{1}{3}u_{2}(e,\theta_{2})=8/3$
from truth-telling and an expected payoff of
$\frac{1}{3}u_{2}(b,\theta_{2})+\frac{1}{3}u_{2}(c,\theta_{2})+\frac{1}{3}u_{2}(d,\theta_{2})=10/3$
from deceiving to $\theta_{2}^{\prime}$, so the best response of type $\theta_{2}$ is to deceive. Similarly type $\theta_{2}^{\prime}$ will deceive to $\theta_{2}.$ Thus $(\theta_{2}^{\prime},\theta_{2})$ is agent 2’s best response to $(\theta_{1}^{\prime},\theta_{1}).$ Similarly, $(\theta_{1}^{\prime},\theta_{1})$ is agent 1’s best response to $(\theta_{2}^{\prime},\theta_{2})$. It follows that $\alpha=((\theta_{1}^{\prime},\theta_{1}),(\theta_{2}^{\prime},\theta_{2}),(\theta_{3},\theta_{3}^{\prime}))$ is an undesired equilibrium in the direct mechanism.
It remains to show that $\alpha$ violates Bayesian monotonicity, implying that agents 1 and 2 will always have an incentive to simultaneously deceive in any incentive-compatible mechanism. Recall that $F$ satisfies Bayesian monotonicity iff for all deceptions $\alpha$ such that $F\circ\alpha\neq F$ there exist $i\in I$ and a function $y:\Theta_{-i}\to A$ such that
for all $\theta_{i}\in\Theta_{i}$ and
for some $\theta_{i}^{\prime}\in\Theta_{i}.$
It needs to be shown that for no agent $i$ does there exist an alternate SCF $y$ with the desired properties. First consider agent 1. In order to define a $y:\Theta_{-1}\to A$ such that
$\mathbb{E}(u_{1}(y(\alpha_{-1}(\theta_{-1})),\theta_{1})\mid\theta_{1})>\mathbb{E}(u_{1}(F(\alpha(\theta_{-1},\theta_{1})),\theta_{1})\mid\theta_{1})$
it must be that (1) $y(\theta_{2},\cdot)=e$, $y(\theta_{2}^{\prime},\cdot)\succeq_{\theta_{1}}d$, and at least one of $y(\theta_{2}^{\prime},\theta_{3})\succeq_{\theta_{1}}b,c$ or $y(\theta_{2}^{\prime},\theta_{3}^{\prime})\succeq_{\theta_{1}}b,c$ holds, or (2) the following inequalities hold.
$u_{1}(y(\theta_{2},\theta_{3}),\theta_{1})+u_{1}(y(\theta_{2},\theta_{3}^{\prime}),\theta_{1})<11$
@@ 4.2 Bayesian monotonicity
$u_{1}(y(\theta_{2}^{\prime},\theta_{3}),\theta_{1})+u_{1}(y(\theta_{2}^{\prime},\theta_{3}^{\prime}),\theta_{1})>4.$
If (1) applies, compute that
$\mathbb{E}(u_{1}(y(\theta_{-1}),\theta_{1})\mid\theta_{1})\geq\frac{21}{6}>\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1})\mid\theta_{1})=\frac{10}{3},$
which is not a valid alternate SCF $y.$ Now, suppose (2) applies. Let
$A=u_{1}(y(\theta_{2},\theta_{3}),\theta_{1})+u_{1}(y(\theta_{2},\theta_{3}^{\prime}),\theta_{1}),$
@@
$B=u_{1}(y(\theta_{2}^{\prime},\theta_{3}),\theta_{1})+u_{1}(y(\theta_{2}^{\prime},\theta_{3}^{\prime}),\theta_{1}).$
Then case (2) says that $A<11$ and $B>4.$ Given the utility specifications for agent 1, it follows that $A\leq 9.5$ and $B\geq 5.$ Observe that
$\mathbb{E}(u_{1}(y(\alpha_{-1}(\theta_{-1})),\theta_{1})\mid\theta_{1})>\mathbb{E}(u_{1}(F(\alpha(\theta_{-1},\theta_{1})),\theta_{1})\mid\theta_{1})\Rightarrow 2A+B>26,$

### v4

@@ 3 Partial implementation and the direct mechanism
In this section we show that providing additional information to agents about the types of other agents is not helpful for the planner interested only in partial implementation and set up our full implementation analysis. We want to show that if $F$ is not implementable with respect to an uninformative signal structure, then it is not implementable with respect to any signal structure $\{\sigma_{i}\}_{i\in I}$. For the remainder of this paper, write $\phi_{i}$ to mean an uninformative signal about $\Theta_{-i}$ for each $i.$
@@
The law of total expectations implies that (3) is just a weighted average of the equations  (3) for $1\leq\ell\leq k.$ This implies the following proposition.
@@
It is well-known that the direct mechanism generally has undesired equilibria. The mechanism exhibited by Jackson (1991) to fully implement an SCF $F$ allows each agent to report, among other information, an alternate SCF that can be used to signal a deception to the planner, provided $F$ is Bayesian monotonic. We introduce the following notion to keep track of which agents have an incentive to alert the planner of a given deception.
@@
The inequalities are exactly the ones from the definition of Bayesian monotonicity. Since a deception is equivalent to a pure-strategy profile in the direct mechanism, the only deceptions $\alpha$ such that $F\circ\alpha\neq F$ that need to be undermined in order to guarantee the full implementation of $F$ are (untruthful) equilibria of the direct mechanism. For any incentive-compatible SCF $F$, write $\mathcal{A}_{F}$ to be the set of deceptions $\alpha$ such that $F\circ\alpha\neq F$ and constitute a pure-strategy equilibrium of the direct mechanism when agents do not have additional information about other agents. Let $\mathcal{A}_{F}^{u}\subset\mathcal{A}_{F}$ be the set of deceptions that can be undermined by some agent $i,$ again when agents have no additional information about other agents. Then we have the following.
@@
Let $F$ be an incentive-compatible SCF. Then $F$ is fully implementable if and only if $\mathcal{A}_{F}^{u}=\mathcal{A}_{F}$.
The lemma follows from the mechanism constructed in the full Bayesian implementation result of Jackson (1991). It allows us to focus our attention on a small and well-defined set of deceptions rather than the unruly space of all possible deceptions suggested by the definition of Bayesian monotonicity.
4 An illustrative example
The goal of this paper is to provide a sufficient condition for a non-fully-implementable SCF to be fully implementable when agents can receive noisy information about other agents’ types. In this section, we illustrate an example that shows how an agent for which truthful reporting is a dominant strategy can use information about other agents to signal deceptions to the planner.
Suppose a firm selling a product is deciding whether to invest more in the product ($I$), maintain the status quo ($S$), or exit the market ($E$). To decide what to do, the board (planner) designs a mechanism between a senior analyst (agent 1), junior analyst (agent 2), the CEO (agent 3), and a compliance officer (agent 4). Each analyst $i$ privately observes a signal in $\{H_{i},L_{i}\}$ regarding whether demand for the product is high ($H_{i}$) or low ($L_{i}$), which constitute’s agent $i$’s private type $\theta_{i}$. The CEO and compliance officer have only one type (or equivalently do not have private information), and the analysts’ types are correlated according to the common prior distribution given by
$\mu(\theta_{1}=H_{1},\theta_{2}=H_{2})=\mu(\theta_{1}=L_{1},\theta_{2}=L_{2})=\frac{1}{3},$
@@ 4 An illustrative example
$\mu(\theta_{1}=H_{1},\theta_{2}=L_{2})=\mu(\theta_{1}=L_{1},\theta_{2}=H_{2})=\frac{1}{6}.$
An SCF maps the pair of analysts’ type to a triple $(y,t_{1},t_{2})$ consisting of a firm action $b\in\{I,S,E\}$ and non-negative transfers $t_{1},t_{2}$ received by the analysts. The firm has 2 indivisible units of money that can be paid out to analysts, so the outcome set is
$A=\{I,S,E\}\times\{(t_{1},t_{2})\in\mathbb{N}^{2}\mid t_{1}+t_{2}\leq 2\}.$
All agents have type-independent preferences, so we omit types from the utility functions in this example. Each analyst’s utility is equal to his transfer. The CEO’s utility function is given by
$u_{3}(b,t_{1},t_{2})=9\mathbbm{1}(b=I)+6\mathbbm{1}(b=S)+3\mathbbm{1}(b=E)-t_{1}-t_{2},$
reflecting that she always prefers investing to keeping status quo to exiting, and within each action prefers to give spend less money on the analysts. Finally, the compliance officer’s utility is
$u_{4}(b,t_{1},t_{2})=2\mathbbm{1}(b=S)-\mathbbm{1}(b=I),$
reflecting a preference for keeping status quo to exiting to investing. The board’s objective is given by the SCF
$F(\theta_{1},\theta_{2})=\begin{cases}(I,1,1)&\text{if }(\theta_{1},\theta_{2})=(H_{1},H_{2}),\\[4.0pt] (S,0,0)&\text{if }(\theta_{1},\theta_{2})=(H_{1},L_{2}),\\[4.0pt] (E,0,0)&\text{if }(\theta_{1},\theta_{2})=(L_{1},H_{2}),\\[4.0pt] (E,1,1)&\text{if }(\theta_{1},\theta_{2})=(L_{1},L_{2}).\end{cases}$
In other words, the board wants to pay the analysts $(1,1)$ only when they deliver the same recommendation. The firm values the senior analyst’s opinion more and is risk-averse, exiting whenever the senior analyst reports low demand and maintains the status quo if the senior analyst reports high demand but the junior analyst disagrees. The board only wants to invest if both analysts agree the demand is high. In the remainder of this section, we show that $F$ is not fully implementable under an uninformative signal structure and construct a signal that can be provided to the CEO about the analysts’ that allows the board to fully implement $F.$
First consider the implementability of $F$ without additional signals. As noted in Lemma 1, $F$ is fully implementable if and only if every equilibrium $\alpha$ of the direct mechanism such that $F\circ\alpha\neq F$ can be undermined by some agent. The direct mechanism, which allows only the analysts to have nontrivial strategy spaces in this case, has exactly four pure-strategy Bayes Nash equilibria. Note that neither the CEO nor the compliance officer can affect the outcome in the direct mechanism, since they have only one type and truth-telling is trivially their dominant strategy.
  1. 1.
    Truth-telling: Analysts report truthfully, so $s_{i}(H_{i})=H_{i}$ and $s_{i}(L_{i})=L_{i}$ for $i=1,2.$
  2. 2.
    Always-$H$: $s_{i}(H_{i})=s_{i}(L_{i})=H_{i}$.
  3. 3.
    Always-$L$: $s_{i}(H_{i})=s_{i}(L_{i})=L_{i}$.
  4. 4.
    Always-lie: $s_{i}(H_{i})=L_{i}$ and $s_{i}(L_{i})=H_{i}$.
Proposition 2.
The compliance officer can undermine both the always-$H$ and always-$L$ equilibria, but no agent can undermine the always-lie equilibrium.
Proof.
Recall that an agent can undermine a deception (undesired equilibrium of the direct mechanism) if she can propose an alternate SCF that weakly harms her payoff, regardless of her own type, under truth-telling but strictly benefits her payoff under the deception. If the analysts are playing the always-$H$ deception, then the compliance officer can propose the following alternate SCF:
$y(\theta_{1},\theta_{2})=\begin{cases}(S,0,0)&\text{if }(\theta_{1},\theta_{2})=(H_{1},H_{2}),\\[4.0pt] (I,1,1)&\text{otherwise}.\end{cases}$
It is readily verified that $y$ gives the compliance officer the same expected utility as $F$ if the analysts were truthfully reporting but strictly benefits her if the analysts are playing always-$H$, implying that the compliance officer can undermine the always-$H$ deception. A similar argument shows the compliance officer can undermine the always-$L$ deception. To show $F$ is not fully implementable, it remains to show no agent can undermine the always-lie deception. To see this, first consider the senior analyst’s perspective. Any alternate SCF $y$ he can propose must only depend on $\theta_{2}$, so let $y(H_{2})=a_{1}$ and $y(L_{2})=a_{2}.$ If the senior analyst can undermine always-lie, $y$ must weakly harm his utility under truth-telling, so $a_{1},a_{2}$ must be chosen such that
$\frac{2}{3}u_{1}(a_{1})+\frac{1}{3}u_{1}(a_{2})\leq\frac{2}{3}$
@@
$\frac{1}{3}u_{1}(a_{1})+\frac{2}{3}u_{1}(a_{2})\leq\frac{2}{3}.$
But for the senior analyst to have an incentive to propose $y$ if the analysts are playing always-lie, $y$ must satisfy either
$\frac{2}{3}u_{1}(a_{1})+\frac{1}{3}u_{1}(a_{2})>\frac{2}{3}$
or
$\frac{1}{3}u_{1}(a_{1})+\frac{2}{3}u_{1}(a_{2})>\frac{2}{3},$
contradicting the previous requirement. Thus the senior analyst cannot undermine the always-lie deception. A similar argument applies to the junior analyst. Finally, both the CEO and compliance officer are always indifferent between truth-telling and always-lie under any SCF, so neither agent can undermine always-lie. [^10] This indifference is not necessary in this example or in general but simplifies the exposition.
Therefore, $F$ is not fully implementable under an uninformative signal structure.
Proposition 3.
$F$ is fully implementable with respect to $\{\sigma_{i}\}$, where $\sigma_{1},\sigma_{2},$ and $\sigma_{4}$ are uninformative and $\sigma_{3}$ has support in $\{0,1\}$ and satisfies
$Pr(\sigma_{3}=0\mid\theta_{-3}=(H_{1},H_{2}))=Pr(\sigma_{3}=0\mid\theta_{-3}=(L_{1},H_{2}))=\tau$